Calculate the effective interest rate for each following normal interest rates




a) 3% compounded quarterly. Round to two decimal point




b) 3% compunded month. Round to two decimal point

1
Expert's answer
2022-03-08T07:39:02-0500

The effective interest rate is given as

"(1+\frac{i}{n})^{n}-1"

a) at 3% compounded quarterly

n= 4


"=(1+\frac{0.03}{4})^{4}-1"

"=3.03\%"


b) at 3% compounded monthly

n= 12

"=(1+\frac{0.03}{12})^{12}-1"

"=3.04\%"